Home‎ > ‎Crew 133‎ > ‎

114-CommanderReport

Commander Report
01/14/2014
Paula Crock
 
Psychological flexibility.  There’s probably a more precise phrase for what I mean but that one will
do.  It’s part of the astronaut selection process and today I saw a great example from the crew.
The plan was for Gartrelle, Howell, Diaz-Rubin and Komorowski to take the ATVs north to check some
strata points and collect helmet temperature data.  I had started the ATVs to get them warmed up,
then helped the crew get suited up but by the time they got outside two of the ATVs were not
operating well.  Without any hesitation the crew suggested they head off for a walking EVA and reach
whatever points they could on foot to complete the mission as much as possible.  It was a tough walk
for them in muddy conditions and a warm day but they persevered and completed much of the mission
objectives.
 
Indoors we had a good day too.  The lettuce seedlings that Gartrelle and Howell started last week
were finally moved into their permanent homes in the hydroponics pods this morning.  The hydroponics
system seems to be off to a good start with minimal leaking and healthy transplants this evening.

=====

CommanderCheckIn Report
01/14/2014
Paula Crock
Crew Physical Health:
All are well.
  
Eva Depart Time:
11:05
Eva Return Time:
13:46
Narrative Of Field Mission Results:
Radio comms study data and helmet temperature data collected
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Hab Maint:
ATV#1 lost some oil through loose dipstick during warmup idle, ATV#2 idling high (ATV EVA changed
to walking EVA due to ATV issues)
Second radio astro receiver sent from Detterline has been fixed
Water tank replaced with full tank
 
Rep Sched:
Commander Check-In
Engineering Report
Science Report (GreenHab)
Science Report (Radio Astro)
Science Report (Astro)
EVA Plan
EVA Report
Daily Photos
Journalist Report
Commander Report
Tomorrow Plan:
All crew in-Hab for final project push and mid-week cleaning
Inventory:
N/A
Support Req:
Continue to await word on whether there are replacement bulbs/ballasts on station for the lab
lights
Miscelleneous:
Lettuce plants were transplanted to hydroponics pods today